The Latest from the LHC: hitting the target luminosity for 2010!

Thanks to a significant increase in the number of bunches in each beam, the 2010 target peak luminosity of 1032 cm-2 s-1 was reached on 14 October 2010, with further progress made in the following days. Soon, the attention of the LHC operators will turn to operation with lead ions throughout November.

In the last two weeks the number of bunches injected in each beam has steadily increased to reach 312, of which 295 collide in points 1 (ATLAS experiment), 5 (CMS experiment) and 8 (LHCb experiment). This has allowed the operators to reach a luminosity of 1.48x1032 cm-2 s-1, comfortably exceeding the target for 2010.

With the present number of bunches, there are over 3.5x1013 protons per beam and around 20MJ of stored energy per beam. Since 4 October, when 204 bunches per beam were injected into the LHC, some intensity-related effects have started to be observed, notably, a significant rise in the ATLAS background. This is linked to an increase in pressure in the beam vacuum about 60m either side of the interaction point. Investigations have shown that this phenomenon actually started when bunch train operation began on 22 September and that it gets worse as the number of bunch trains increases. It is also linked to the bunch structure in the trains and is only present with two beams in the machine. This interesting effect is certainly not a show-stopper, and lessens with time as the circulating beams clean the surface of the vacuum chamber in this location.

On Saturday 9 October experts installed some additional solenoids in the regions where the most significant background had been observed. By powering them when a pressure rise was observed in the vacuum pipe, they were able to demonstrate that turning the solenoid on reduced the pressure rise. This seems to indicate that much – but not all – of the pressure rise is due to electron cloud effects, that is, synchrotron radiation from the beam when it hits the beam screen and knocks out electrons, which can, in turn, hit the wall and knock out more electrons.

Over the last week, another problem has come to light. Starting on Friday 8 October, increased losses have been observed when injecting Beam 1 into point 2 of the LHC. It has been possible to improve the situation by re-steering the beam through the injection channel, but then the problem became worse on Saturday 16 October. Again it was possible to re-steer through the injection channel, but now there is almost no margin. To investigate the source of the problem, radiation surveys and X-ray imaging have been made, clearly showing an aperture restriction at the transition between two magnetic septa in the injection channel, thought to be due to a non-conformity in the mounting of the interconnection. Since even higher intensity running is foreseen in 2010, it has been decided to interrupt machine operation and make the necessary repairs. This action needs a few days to complete, hence the technical stop foreseen for 1 to 4 November has been advanced to 19 to 22 October, thereby minimizing the impact on the physics program. Proton operation will continue from Friday 22 October to Friday 5 November, when attention will be switched to operation with lead ions for the remainder of the 2010 run.



by CERN Bulletin